OpenStudy (anonymous):

Solve for t: e^ln(0.6)t = 0.9

OpenStudy (bibby):

e^x and ln(x) are inverse operations. As such e^ln(t) = t e^ln(0.6)t = 0.9 0.6t = 0.9

OpenStudy (anonymous):

Thanks! That's what I thought it was, but wasn't sure. So, 1.5 is the answer?

OpenStudy (bibby):

yep. 0.9/0.6 = 1.5

OpenStudy (anonymous):

Thank you!!!